OpenStudy (anonymous):

somebody please help me- general admission tickets to the fair cost $3.50 per person. ride passes cost an additional $5.50 per person. parking costs $6 for the family. the total costs for ride passes and parking was $51. how many people in the family attended the fair?

OpenStudy (anonymous):

let x=number of people purchasing tickets cost per person=general+ride=3.50+5.50=9 per person then total cost=cost per person times number of persons + parking 9x+6=51 9x=45 x=6 people in the family (if I am interpreting the problem correctly)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

ur welcome :})